# multi-AUV; task assigment; kinematic constraints; Dubins path; SOM Neural Network

### v3

@@ 0 Introduction
The market mechanism algorithm is one the most common method to solve the multitask assignment operations Oh et al. (2015). Another commonly used task assignment algorithm is ant colony algorithm which mainly studies the autonomous task assignment of multi-robot systems in dynamic environments Cao et al. (2013). These are also other task assignment methods, such as that with intelligent heuristic algorithm handling water flow influence Zhu and Yang (2024), while kinematic path planning not considered for most of them. As for neural networks, literature Liu et al. (2011) further proposed a task allocation algorithm suitable for large-scale multi-robot clusters. The self-organizing neural network algorithm can also be applied to multi-task assignment. The SOM algorithm was proposed by a Finnish scholar, Teuvo Kohonen, in the 1980s Kohonen (1982); Zhao and Guo (2020). Literature Zhu and Yang (2006); Zhu et al. (2012) applied the self-organizing neural network to the task assignment and distribution of multi-mobile robot systems. path planning. Using the SOM neural network algorithm, dynamic task assignments for multi-AUV systems in two-dimensional ocean environments are implemented. Then, the multi-AUV multi-target allocation strategy of self-organizing map (SOM) neural network is further extended and applied to the three-dimensional marine environment Li and Zhu (2018).
@@
As shown in Figure 1, a typical AUV has 1 to 2 thrusters at the tail, relying on rudder and fins to control the moving direction. This structure determines that most AUVs are underactuated. When an AUV moves, it turns horizontally and tilts vertically, i.e., yaw and pitch. The yaw and pitch angles are limited due to mechanical structure and fluid dynamics. In 3D workspace, pitch angle $\gamma\leq\gamma_{max}$, and yaw angle’s derivative or the angular velocity on X-Y plane $\dot{\varphi}$ is also limited, which results the turning radius $r\geq R$, where $R$ is the minimum turning radius under the angular velocity of $max(\dot{\varphi})$.
It should be noted that when the force of the thruster is set very high, the turning radius will also be large at the maximum rudder angle while we are discussing the case of normal thrust.
Considering this, task assignment and path planning of multiple AUV systems under kinematic constraints in a two-dimensional environment are studied in this paper, taking into account static obstacles and other interference factors. The main innovation is to combine the task assignment of SOM neural networks with Dubins path planning under kinematic constraints, making the method more applicable to real-world applications. By combining the Dubins Path algorithm with the improved self-organizing mapping (SOM) neural network algorithm, this paper proposes a new task assignment and path planning algorithm, which effectively solves the multi-task assignment problem of the multi-AUV system in the actual environment. The mathematical model and algorithms are established on MATLAB, and then tested in Python on a Raspberry Pi.
@@ 2 Task assignment algorithm under kinematic constraints
In this section, self-organizing map (SOM) neural network method with Dubins path generator is used to deal with task allocation and path planning problems for the multi-AUV system. The control stability and accuracy of the system model are compensated by strengthening the optimal global exploration and local exploitation ability. At the same time, the improved optimization method can realize the controller’s online parameter adjustment to meet the AUVs’ control requirements in the experimental simulation environment.
2.1 Application of SOM in Multi-AUV Task Assignment
@@
$D_{kjl}=\left|\mathbf{T}_{l}-\mathbf{R}_{jk}\right|(1+V)$ (6)
and
@@
The flowchart of the algorithm is shown in Figures 4. After initializing the SOM neural network, the positions of the target points are sequentially input into the network. For a given input target point during the iteration process, it can be summarized as a three-step calculation process. The first step is to select the winning neuron; The second step is to determine the neighborhood of the winning neuron; The third step is to correct the weight vectors of the winning neuron and its neighboring neurons. After the specified AUV reaches all target points, task allocation is completed.
@@
Figure 4: The SOM algorithm flow chart.
@@ 2.2 Dubins path planning algorithm for AUV kinematic constrains
Unlike underwater robots such as ROVs, most torpedo like AUVs and underwater gliders are underactuated with tail thrusters to only moving forward, steering depend on fins and rudders. It is difficult to achieve precise underwater motion control, which poses high requirements for path planning. Fortunately, Dubin’s Car was introduced into the literature by Lester Dubins, a famous mathematician and statistician, in a paper published in 1957. The cars essentially have only 3 controls: “turn left at maximum”, “turn right at maximum”, and “go straight”. All the paths traced out by the Dubin’s car are combinations of these three controls. Dubin’s car is very similar to AUVs’ kinematics. Having precise motion control to solve dynamic problems is also very useful, which is another quite challenging field being studied and will be involved in our future work. Here, we adopt the Dubins method to deal with three-dimensional motion problems.
Literally, the Dubins curve is the shortest path connecting two points, while satisfying curvature constraints and specified tangent lines (entry direction) at the beginning and end, and limiting the target to only move forward. As described in prior works, the Dubins curve can be represented as a combination of three basic movements, as listed in Table 1. The Dubins curve provides a sufficient set of paths, which includes the optimal path. The shortest path is only selected from 6 curves in the set of LRL LSL LSR RLR RSR RSL.
@@ C Circular arc Na
where $z_{0}$ and $z_{1}$ are the Z coordinates of the starting and ending points. Detailed procedure has been introduced in existing papers such as Cai et al. (2017); Vana et al. (2020) and will not be repeated here.

### v4

@@ 0 Introduction
The market mechanism algorithm is one the most common method to solve the multitask assignment operations Oh et al. (2015). Another commonly used task assignment algorithm is ant colony algorithm which mainly studies the autonomous task assignment of multi-robot systems in dynamic environments Cao et al. (2013). These are also other task assignment methods, such as that with intelligent heuristic algorithm handling water flow influence Zhu and Yang (2024), while kinematic path planning not considered for most of them. As for neural networks, literature Liu et al. (2011) further proposed a task allocation algorithm suitable for large-scale multi-robot clusters. The self-organizing neural network algorithm can also be applied to multi-task assignment. The SOM algorithm was proposed by a Finnish scholar, Teuvo Kohonen, in the 1980s Kohonen (1982); Zhao and Guo (2020). Literature Zhu and Yang (2006); Zhu et al. (2012) applied the self-organizing neural network to the task assignment and distribution of multi-mobile robot systems. path planning. Using the SOM neural network algorithm, dynamic task assignments for multi-AUV systems in two-dimensional ocean environments are implemented. Then, the multi-AUV multi-target allocation strategy of self-organizing map (SOM) neural network is further extended and applied to the three-dimensional marine environment Li and Zhu (2018). Event-triggered adaptive neural network tracking control for uncertain systems was proposed in Liu et al. (2024), where the event-triggering mechanism can reduce the update rate of input signals and avoid the Zeno behavior.
@@
As shown in Figure 1, a typical AUV has 1-2 thrusters at the tail, relying on rudder and fins to control the moving direction. This structure determines that most AUVs are underactuated. When an AUV moves, it turns horizontally and tilts vertically, i.e., yaw and pitch. The yaw and pitch angles are limited due to mechanical structure and fluid dynamics. In 3D workspace, pitch angle $\gamma\leqslant\gamma_{max}$, and yaw angle’s derivative or the angular velocity on X-Y plane $\dot{\varphi}$ is also limited, which results the turning radius $r\geqslant R$, where $R$ is the minimum turning radius under the angular velocity $max(\dot{\varphi})$. It should be noted that when the force of the thruster is set very high, the turning radius will also be large at the maximum rudder angle, while we are discussing the case of normal thrust. Considering this, task assignment and path planning of multiple AUV systems under kinematic constraints in a two-dimensional environment are studied in this paper, taking into account static obstacles and other interference factors. The main innovation is to combine the task assignment of SOM neural networks with Dubins path planning under kinematic constraints, making the method more applicable to real-world applications. By combining the Dubins Path algorithm with the improved self-organizing mapping (SOM) neural network algorithm, this paper proposes a new task assignment and path planning algorithm, which effectively solves the multi-task assignment problem of the multi-AUV system in the actual environment. The mathematical model and algorithms are established on MATLAB, and then tested in Python on a Raspberry Pi.
@@ 2 Task assignment algorithm under kinematic constraints
In this section, self-organizing map (SOM) neural network method with Dubins path generator is used to deal with task allocation and path planning problems for the multi-AUV system. The control stability and accuracy of the system model are compensated by strengthening the optimal global exploration and local exploitation ability. At the same time, an event-triggered neural network evolution strategy is developed to decrease the update frequency of the control signal. The improved optimization method can realize the controller’s online parameter adjustment to meet the AUVs’ control requirements in the experimental simulation environment.
2.1 Application of Event-triggered SOM in Multi-AUV Task Assignment
@@
$\left.D_{kjl}=\begin{cases}\left|\mathbf{T}_{l}-\mathbf{R}_{jk}\right|\left(1+V\right),&P_{j}<S_{\max}\\ \infty,&P_{j}\geq S_{\max}\end{cases}\right.$ (6)
where $S_{\max}$ is the maximal distance that a single AUV can travel, and
@@
At the same time, if there exists an obstacle at the position $ob=(x_{o},y_{o})$, the obstacle weight is computed based on the Euclidean distance between it and the neural network weights by
$ob_{w}=min(\left|\mathbf{ob}-\mathbf{R}_{jk}\right|)$ (11)
where $\mathbf{ob}$ is the expansion of $ob$ to a vector of equal length to $\mathbf{R}_{jk}(\mathrm{j}=1,2,\ldots,\mathrm{J};\mathrm{k}=1,2,\ldots,\mathrm{K})$. By comparing $ob_{w}$ with the defined safety distance $d_{s}afety$, the impact of obstacles can be determined.
The flowchart of the algorithm is shown in Figures 4. After initializing the SOM neural network, the positions of the target points are sequentially input into the network. For a given input target point during the iteration process, it can be summarized as a three-step calculation process. The first step is to select the winning neuron; the second step is to determine the neighborhood of the winning neuron; the third step is to correct the weight vectors of the winning neuron and its neighboring neurons. Event triggering occurs in three situations, namely obstacles on the path, load exceeding the upper limit, or inability to plan the Dubins path from the AUV to the pre-allocated target point. The event triggered control functions $u$ is given by (12), where $u_{1}$ is the workload related function by (13) and $u_{2}$ is the obstacles related function by (14). The event triggered re-assignment will be implemented in the SOM algorithm flow chart if $u=0$. After all target points are reached by one specified AUV, task allocation is completed.
$u=u_{1}\&u_{2}$ (12)
$u_{1}=\begin{cases}0,&D_{kjl}=\infty\\ 1,&others\end{cases}$ (13)
$u_{2}=\begin{cases}0,&ob_{w}>d_{safety}\\ 1,&ob_{w}\leq d_{safety}\end{cases}$ (14)
@@
Figure 4: The event-triggered SOM algorithm flow chart.
@@ 2.2 Dubins path planning algorithm for AUV kinematic constrains
Unlike underwater robots such as ROVs, most torpedo like AUVs and underwater gliders are underactuated with tail thrusters to only moving forward, steering depend on fins and rudders. It is difficult to achieve proper path planning. Fortunately, Dubin’s Car was introduced into the literature by Lester Dubins, a famous mathematician and statistician, in a paper published in 1957. The cars essentially have only 3 controls: “turn left at maximum”, “turn right at maximum”, and “go straight”. All the paths traced out by the Dubin’s car are combinations of these three controls. Dubin’s car is very similar to AUVs’ kinematics. Assuming that the AUV has an initial velocity and dynamic characteristics as mentioned above, the path can be planned based on the Dubins car method from 2D to 3D workspace. Precise motion control to dynamic control problem is also very useful, which is another quite challenging field being studied. For example, command filter with adaptive control was proposed to solve the model uncertainties and input saturation issues Liu et al. (2023). These theories will be involved in our future work. To minimize the traveling time and energy consumption of AUVs, path planning based on kinematic is necessary. Here, we adopt the Dubins method to deal with the three-dimensional motion planning problem.
Literally, the Dubins curve is the shortest path connecting two points with initial directions, while satisfying curvature constraints and specified tangent lines (entry direction) at the beginning and end, and limiting the target to only turn forward. As described in prior works, the Dubins curve can be represented as a combination of three basic movements, as listed in Table 1. The Dubins curve provides a sufficient set of paths, which includes the optimal path. The shortest path is only selected from 6 curves in the set of LRL LSL LSR RLR RSR RSL.
@@ C Circular arc Na
where $z_{0}$ and $z_{1}$ are the Z coordinates of the starting and ending points. Detailed procedure has been introduced in existing papers such as Cai et al. (2017); Vana et al. (2020) and will not be repeated here. If obstacles appear in the path, we must re-assign the tasks.